OpenStudy (anonymous):

A student in an algebra course has test scores of 40,55 and 72. What score in the 4th test will give the student an average of 60?

OpenStudy (gw2011):

(40+55+72+x)/4=60 40+55+72+x=240 x=240-40-55-72 x=73